At a theme park, Setha went on twice as many rides as Chenda, who went on three more rides than Dara. Together, they went on 25 rides. How many rides did Dara go on?
Found 2 solutions by solver91311, ikleyn:
Answer by solver91311(24713)



Let represent the answer you are asked for.



Solve for

John

My calculator said it, I believe it, that settles it


I > Ø

Answer by ikleyn(52781)

.

Dara   = x
Chenda = (x+3)
Setha  = 2*(x+3)


    x + (x+3) + 2*(x+3) = 25


etc.